# What Is a Good Caption? A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

### v2

@@ Abstract
Visual captioning benchmarks have become outdated with the emergence of modern multimodal large language models (MLLMs), as the brief ground-truth sentences and traditional metrics fail to assess detailed captions effectively. While recent benchmarks attempt to address this by focusing on keyword extraction or object-centric evaluation, they remain limited to vague-view or object-view analyses and incomplete visual element coverage. In this paper, we introduce CAPability, a comprehensive multi-view benchmark for evaluating visual captioning across 12 dimensions spanning six critical views. We curate nearly 11K human-annotated images and videos with visual element annotations to evaluate the generated captions. CAPability stably assesses both the correctness and thoroughness of captions using F1-score. By converting annotations to QA pairs, we further introduce a heuristic metric, know but cannot tell ($K\bar{T}$), indicating a significant performance gap between QA and caption capabilities. Our work provides the first holistic analysis of MLLMs’ captioning abilities, as we identify their strengths and weaknesses across various dimensions, guiding future research to enhance specific aspects of capabilities.
Figure 1: An example of image caption (left) and video caption (right). By analyzing the components of captions, we conclude 12 dimensions (9 static dimensions and 4 dynamic dimensions with object number shares on both static and dynamic), which all contribute to a detailed and comprehensive caption. The static dimensions are shared in both images and videos. For video data, there are additional dynamic dimensions as they need to be judged with temporal relations.
* Interns at Alibaba Group
Corresponding Author
@@ 1 Introduction
Visual captioning, which translates visual content into textual descriptions, is a fundamental task for both image and video understanding [15, 41], and forms a significant basis for image and video generation [29, 34]. To assess the capabilities of this task, researchers established several visual caption benchmarks in earlier years [8, 2, 40, 39].
However, with the rapid development of recent MLLMs [21, 47, 26, 19, 30, 9, 24, 44, 32], these traditional benchmarks have become outdated. This can be attributed to two main reasons: 1) The ground truths of traditional benchmarks often contain short sentences, missing many details. In contrast, recent MLLMs can produce much more detailed and fine-grained captions than the ground truths. 2) Traditional benchmarks use N-gram-based metrics, such as BLEU [28] and CIDER [33], to directly compare the similarity between generated captions and ground-truth sentences, making evaluations unreliable due to their high sensitivity to sentence style.
Figure 2: The development of visual caption benchmarks. Many works compare the ground-truth with generated sentences, which is vague. CompreCap [25] uses a scene graph to evaluate only object-related information in caption. Our CAPability considers multiple views and conducts comprehensive evaluation.
Table 1: Views of our designed dimensions. We can treat a caption from the listed six views, and then split each of them into several dimensions.
@@
Recently, new visual caption benchmarks have been introduced to update the outdated ones. As illustrated in Fig. 2, Dream-1K [35], DetailCaps [13] and VDC [7] extract keywords from both generated and reference captions, (e.g., objects in DetailCaps, events in Dream-1K, objects, background, and camera information in VDC), and then compare the extracted information to score the caption, as opposed to earlier methods that directly compared sentences. We name all these methods vague-view evaluation as their evaluations still depend on the level of detail and accuracy of the ground-truth caption, which can suffer from human bias and cumulative errors from repeated extraction and comparison by LLMs. CompreCap [25] extracts object-related annotations from images (e.g., scene graph) without ground-truth captions, thereby focusing on evaluating the object description capabilities of modern image MLLMs. We refer to this as object-view evaluation, as it drops entire sentences as ground truth, and evaluates captions based on object representation. Compared to traditional benchmarks, all these newly introduced approaches aim to provide more precise ground truths and evaluation methods, enhancing the reliability and interpretability of benchmarking.
@@
However, the evaluation of these benchmarks remains incomplete as they focus on a single aspect of captions with limited visual elements, inadequately covering the full caption scope. For instance, they often overlook aspects like scene, text, and style. In this paper, we conduct a multi-view evaluation and introduce a new comprehensive visual caption benchmark, CAPability, with 6 views (i.e., object, global, text, camera, temporal, and knowledge) and 12 dimensions. This approach uses complete visual elements rather than caption sentences as annotations to evaluate both correctness and thoroughness for each dimension. The relationship between 6 views and 12 dimensions is listed in Tab. 1, and the designed dimensions are illustrated in Fig. 1. We believe these components contribute to a complete caption, as lacking any of them may align the caption with different visual content. There are 9 static and 4 dynamic dimensions, with object number encompassing both static and dynamic aspects. Dynamic dimensions apply to both images and videos, while static ones are exclusive to video. We collect and manually annotate 11K images and videos for CAPability, providing sufficient samples.
Figure 3: F1-score comparison of SOTA MLLMs on our CAPability. Gemini-1.5-pro [30] achieves the best.
In addition to multi-view annotation, we also focus on improving the evaluation of captions. While most methods assess only the correctness, we argue that considering both correctness and thoroughness of visual elements provides a more comprehensive evaluation for visual captioning. Therefore, we conduct comprehensive experiments using F1-score as our main metric, supplemented by two other heuristic metrics: hit rate and the know but cannot tell ($K\bar{T}$) metric. The F1-score combines precision and recall to reflect both correctness and thoroughness. The hit rate measures dimension-referential thoroughness, while $K\bar{T}$ indicates when a model can answer related questions correctly but fails to convey the same information in the caption automatically. These metrics provide a robust framework for evaluating both correctness and thoroughness in MLLMs. To our knowledge, we are the first to heuristically highlight the gap in correctness and thoroughness capabilities of MLLMs across multiple views, guiding researchers to enhance these capabilities across dimensions. Representative results are shown in Fig. 3, leading to the following conclusions: 1) Gemini-1.5-pro [30] performs the best across many dimensions, followed by GPT-4o [27]. 2) Gemini-1.5-pro demonstrates strong object-counting abilities, while GPT-4o excels in identifying camera angles. 3) All models still struggle with dimensions like object numbers, camera angle, camera movement, character identification, and action. We hope our findings guide researchers to focus on improving these abilities in caption tasks. Our main contributions are listed as follows:
  • We emphasize that a good caption should be evaluated for both correctness and thoroughness. Accordingly, we report precision, recall, and hit rate, using the F1-score as the main metric to combine correctness and thoroughness.
2 Related Work
Multi-modal large language models. Based on the significant development of Large Language Models (LLMs) among various linguistic tasks [5, 11, 42, 14], many works try to extend the powerful capabilities into multi-modal understanding. By integrating image content into LLMs, Multi-modal Large Language Models (MLLMs) also gain huge achievements [21, 22, 3, 47, 23, 10, 26]. Based on the pre-trained weights from image models, recent MLLMs also expand video understanding capabilities [45, 19, 46, 36, 9, 24, 44, 32]. With rapid development, MLLMs are powerful enough to describe both the image and video content in detail, which makes the traditional benchmarks with short captions outdated. More and more methods even try to produce re-captioned detailed descriptions by more powerful models rather than existing human-annotated short captions to train their model [19, 44]. Therefore, it is urgent to propose a new visual caption benchmark that adapts to modern MLLMs.
@@ 2 Related Work
Table 2: Comparison of our CAPability and other visual caption benchmarks in different aspects. We are the most comprehensive with both image and video data, multi-view annotations, and new thoroughness evaluation methods proposed.
Benchmark Data Type Anno- tations Thoroughness Evaluation
Image Video Recall Hit rate $\bm{K\bar{T}}$
MS-COCO[8] - Sentences - - -
MSRVTT [40] - Sentences - - -
Dream-1K [35] - Sentences single dim - -
VDC [7] - Sentences - - -
DetailCaps [13] - Sentences - - -
CompreCap [25] - Object Info single dim - -
@@
Visual caption benchmarks. Visual Caption is a fundamental task in computer vision. Early visual caption benchmarks, such as MS-COCO [8], NoCaps [2], MSR-VTT [40], and VATEX [39], usually contain a short sentence with limited visual information as the ground truth. They also use metrics like BLEU [28], CIDER [33], and METEOR [4] to calculate the matching score directly between two sentences, which is easily affected by the sentence style. Recently, from the annotation aspect, DetailCaps [13] extracts object-related information from the ground-truth caption, Dream-1K [35] splits the ground truth and candidates sentences into events. VDC [7] also extracts the object, background, and camera information from the video captions by question templates. However, they still rely on the ground-truth caption with human-bias, and require existing LLMs to extract and compare multiple times, thus increasing the cumulative error. CompreCap [25] explores directly annotating the object-related information in image captions, making the benchmarking more interpretable. On the contrary, we are the first time to propose a comprehensive visual caption benchmark covering both image and video data with 6 views and 12 dimensions. For evaluation, most methods only focus on correctness. Dream-1K [35] and CompreCap [25] begin to focus on thoroughness and calculate the recall of events or the object coverage in the segmentation map. However, they still remain incomplete as they only evaluate one dimension and limited metrics. We design various metrics about both correctness and thoroughness, which may be ignored by previous work. We summarize the comparison with other visual caption benchmarks in Tab. 2, and we are the most holistic on all listed aspects.
@@ 3.1 Multiple Dimension Data Annotation
Dimension design. As shown in Tab. 1, we conclude 6 views and split them into 12 dimensions based on the analysis of caption cases. As shown in Fig. 1, we design 9 static dimensions for both video and image, and 4 dynamic dimensions for video, covering most of what makes up a visual caption. We classify dimensions as dynamic or static based on the following principle: descriptions obtainable from a single frame are static, those requiring the entire video are dynamic, which are more related to temporal information. For the object number dimension, the number can be counted statically in an image, and also dynamically in a video, which is more challenging [43]. We also design the annotation type of each dimension as two types: open-ended, and specific categories. Specifically, we define 9 categories for style, 4 categories for camera angle, and 7 categories for camera movement. The specific categories of each dimension can be found in Fig. 6. See Sec. B.1 for details of each dimension.
@@
Data collection. For convenience and problem simplification, we only collect image data for static dimensions and video data for dynamic dimensions. This is based on the common sense that the video understanding capabilities for MLLMs are usually built upon sufficient image understanding capabilities [45, 46, 44, 9]. Since an image or a video cannot cover all these dimensions of information, we directly collect data for each dimension independently, and evaluate each dimension separately as sub-tasks. For static dimensions, we mainly collect images from SA-1B [18], COYO-700M [6], Wukong [16], and Wikipaintings [17], and we also crawl a considerable amount of data from multiple websites by ourselves. We also borrow parts of the image data and annotations from CompreCap [25] for the spatial relation dimension. For dynamic dimensions, we crawl and cut videos for camera movement dimension, borrow videos from Dream-1K [35] for action and event dimensions, and borrow videos from VSI-Bench [43] for the dynamic object number dimension. Fig. 4 shows our data sources for each dimension and their proportion.
@@
Pre-Annotation. The annotations may not be unique for different dimensions. For global-related, camera-related, and knowledge-related views, the annotations tend to be unique as an image only belongs to one kind of scene, style, etc. We directly pre-annotate them by extracting the metadata (e.g., style for images in Wikipaintings [17]), or ask SOTA MLLMs to get a preliminary answer. For object-related, text-related, and temporal-related views, there could be multiple objects, texts, or actions in an image or video. However, it is extremely hard to annotate all objects or actions within an image or a video, as the categories of objects can be divided by almost infinite granularity [31, 37, 38]. Therefore, we do not pursue the most comprehensive annotation possible for each single sample, but randomly select only one object from the visual content, and the same for other dimensions, and reflect the accuracy and thoroughness through the evaluation of a large number of samples. We name this strategy as One Represents All. According to the law of large numbers, the distribution of randomly selection can approximate the expectation of covering different granularities of the entire visual content with a large amount of samples, thus ensuring the unbiased nature of the benchmark. Therefore, the key of this annotation strategy is to keep the selection as random as possible. To avoid humans’ bias on selecting, we ask the three SOTA MLLMs, i.e., GPT-4o [27], Gemini-1.5-pro [30], and Qwen-VL-Max [36] to list all objects and actions at the granularity they deem appropriate in an image or video, ask PaddleOCR [20] to list all texts in an image. We finally use Qwen2.5-Max [42] to merge the results together and randomly select one from the merged list to obtain the pre-annotated results. For further object-related dimensions, e.g., object number, object color, and spatial relation, the object selection follows this strategy, then pre-annotate these attributes by MLLMs.
@@
Data balance. Based on the pre-annotation results for each dimension, we conduct data balance strategy to control the difficulty and diversity. For dimensions with specific categories, we try to make the number of each category similar. For dimensions of open-ended descriptions, we count the frequency of the descriptive words, suppress the long-tail distribution, and keep low-frequency words, ensuring the variety. The final annotation distribution is shown as Fig. 6.
@@
Data filtering. We finally conduct data filtering to drop harmful visual content and re-balance the data since many of them are modified manually. The final distribution of each dimension is shown in Fig. 6. Some benchmark examples are shown in Sec. B.4.
Table 3: The precision, recall, and F1-score of closed-source models and 72B open-source models on all dimensions. The precision represents the accuracy of what the models have described. The recall shows how many visual elements in the image can be described correctly. F1-Score is the harmonic mean of precision and recall. For video inputs, we send the whole video for Gemini, and uniformly sample 50 frames for GPT due to the API limitation of maximum number of images.

### v3

# CAPability: A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

@@ Abstract
Visual captioning benchmarks have become outdated with the emergence of modern multimodal large language models (MLLMs), as the brief ground-truth sentences and traditional metrics fail to assess detailed captions effectively. While recent benchmarks attempt to address this by focusing on keyword extraction or object-centric evaluation, they remain limited to vague-view or object-view analyses and incomplete visual element coverage. In this paper, we introduce CAPability, a comprehensive multi-view benchmark for evaluating visual captioning across 12 dimensions spanning six critical views. We curate nearly 11K human-annotated images and videos with visual element annotations to evaluate the generated captions. CAPability stably assesses both the correctness and thoroughness of captions with precision and hit metrics. By converting annotations to QA pairs, we further introduce a heuristic metric, know but cannot tell ($K\bar{T}$), indicating a significant performance gap between QA and caption capabilities. Our work provides a holistic analysis of MLLMs’ captioning abilities, as we identify their strengths and weaknesses across various dimensions, guiding future research to enhance specific aspects of their capabilities.
Figure 1: An example of image caption (left) and video caption (right) task. By analyzing the components of captions, we conclude 12 dimensions (9 static dimensions and 4 dynamic dimensions with object number shares on both static and dynamic), which all contribute to a detailed and comprehensive caption. The static dimensions are shared in both images and videos. Video data has additional dynamic dimensions that need to be judged with temporal relations.
@@ 1 Introduction
Visual captioning, which translates visual content into textual descriptions, is a fundamental task for both image and video understanding fang2015captions ; xu2015show , and forms a significant basis for image and video generation peebles2023scalable ; wang2023modelscope . To assess the capabilities of this task, researchers established several visual caption benchmarks in earlier years mscoco ; nocaps ; msrvtt ; vatex .
With the rapid development of recent MLLMs llava ; minigpt4 ; gpt4v ; llavaov ; gemini1.5 ; internvl2.5 ; nvila ; videollama3 ; qwen2.5vl , these traditional benchmarks have rapidly become outdated. This can be attributed to two main reasons: 1) The ground truths of traditional benchmarks often contain short sentences, missing many details. In contrast, recent MLLMs can produce much more detailed and fine-grained captions. 2) Traditional benchmarks use N-gram-based metrics (e.g., BLEU papinesi2002bleu , CIDER vedantam2015cider ) to directly compare the similarity between sentences, making evaluations unreliable due to their high sensitivity to sentence style.
Figure 2: The development of visual caption benchmarks. Many works compare the ground-truth with generated sentences, which is vague. CompreCap comprecap uses a scene graph to evaluate only object-related information. Our CAPability considers multiple views with a comprehensive evaluation.
Table 1: Our designed views and more detailed dimensions. We can treat a caption from the listed six views, and then split them into several dimensions.
@@
Recently, new visual caption benchmarks have been introduced to update the outdated ones. As illustrated in Fig. 2, Dream-1K dream1k , DetailCaps detailcaps and VDC auroracap extract keywords from both generated and reference captions, (e.g., objects in DetailCaps, events in Dream-1K, objects, background, and camera information in VDC), and then compare the extracted information to score the caption, as opposed to earlier methods that directly compared sentences. We name all these methods vague-view evaluation as their evaluations still depend on the level of detail and accuracy of the ground-truth caption, which can suffer from human bias and cumulative errors from repeated extraction and comparison by LLMs. CompreCap comprecap extracts object-related annotations from images (e.g., scene graph) without ground-truth captions, thereby focusing on evaluating the object description capabilities of modern image MLLMs. We refer to this as object-view evaluation, as it drops entire sentences as ground truth, and evaluates captions based on object representation. Compared to traditional benchmarks, all these newly introduced approaches aim to provide more precise ground truths and evaluation methods, enhancing the reliability and interpretability of benchmarking.
@@
However, the evaluation of these benchmarks remains incomplete as they focus on a single aspect of captions with limited visual elements, inadequately covering the full caption scope. For instance, they often overlook aspects like scene, text, and style. We argue that a multi-view evaluation for visual caption is essential. In this paper, we introduce a new comprehensive visual caption benchmark, CAPability. Our approach uses complete visual elements rather than caption sentences as annotations to evaluate both correctness and thoroughness for each dimension. The selection of dimensions is motivated by existing visual generation benchmarks geneval ; vbench ; t2vcompbench , which is the inverse task of visual captioning. They usually treat and evaluate the generated visual content from different aspects (e.g., objects, scene, style, camera, motion control). Similarly, we design 6 views and 12 dimensions for CAPability by analyzing several visual captions, as illustrated in Tab. 1 and Fig. 1. We believe these components contribute to a complete caption, as lacking any of them may align the caption with different visual content. There are 9 static and 4 dynamic dimensions, with object number encompassing both static and dynamic aspects. Dynamic dimensions apply to both images and videos, while static ones are exclusive to video. We collect and manually annotate 11K images and videos for CAPability, providing sufficient samples.
Figure 3: Precision and hit comparison of SOTA MLLMs on our CAPability. Models perform more variably on hit metric, which evaluates the thoroughness. GPT-4o performs the best on precision, and Gemini-1.5-pro gemini1.5 performs the best on hit.
@@
Table 2: Comparison of our CAPability and other visual caption benchmarks in different aspects. We are the most comprehensive with image and video data, multi-view annotations, and new thoroughness evaluation methods proposed.
Benchmark Data Type Anno- tations Evaluation
Image Video Thoroughness $\bm{K\bar{T}}$
MS-COCOmscoco - Sentences - -
MSRVTT msrvtt - Sentences - -
Dream-1K dream1k - Sentences single dim -
VDC auroracap - Sentences - -
DetailCaps detailcaps - Sentences - -
CompreCap comprecap - Object Info single dim -
@@
In addition to multi-view annotation, we also focus on improving the evaluation quality of captions. While most methods assess only the correctness, we argue that considering both correctness and thoroughness of visual elements provides a more comprehensive evaluation for visual captioning. Therefore, we conduct comprehensive experiments using both precision and hit as our main metric, supplemented by another heuristic metric: the know but cannot tell ($K\bar{T}$). While precision only evaluates the accuracy of captions, hit considers both accuracy and coverage of captions compared with ground truth, and $K\bar{T}$ indicates the ratio when a model can answer related questions correctly but fails to convey the same information in the caption automatically. These metrics provide a robust framework for evaluating both correctness and thoroughness in MLLMs’ detailed captions. To our knowledge, we are the first to heuristically highlight the gap in correctness and thoroughness capabilities of MLLMs across multiple views, guiding researchers to enhance these capabilities across dimensions. Representative results are shown in Fig. 3, leading to the following conclusions: 1) Models perform more variably on hit, which evaluates the thoroughness and may be ignored by previous research. 2) GPT-4o seems the best on precision. For hit, Gemini-1.5-pro gemini1.5 has a leading advantage in many dimensions, followed by GPT-4o gpt4o . 3) Gemini-1.5-pro demonstrates strong object-counting abilities, while GPT-4o excels in identifying camera angles. 4) All models still struggle with dimensions like object numbers, camera angle, camera movement, character identification, and action. We hope our findings guide researchers to focus on improving these abilities in caption tasks. Our main contributions are listed as follows:
  • We emphasize that captions should be evaluated for both correctness and thoroughness. Accordingly, we report precision and hit to combine correctness and thoroughness.
2 Related Work
Multi-modal large language models. Based on the significant development of Large Language Models (LLMs) among various linguistic tasks gpt3 ; vicuna ; qwen2.5 ; llama3 , many works try to extend the powerful capabilities into multi-modal understanding. By integrating image content into LLMs, Multi-modal Large Language Models (MLLMs) also gain huge achievements llava ; llava1.5 ; qwenvl ; minigpt4 ; llavanext ; internvl ; gpt4v . Based on the pre-trained weights from image models, recent MLLMs also expand video understanding capabilities llavanextvideo ; llavaov ; llavavideo ; qwen2vl ; internvl2.5 ; nvila ; videollama3 ; qwen2.5vl . With rapid development, MLLMs are powerful enough to describe both the image and video content in detail, which makes the traditional benchmarks with short captions outdated. More and more methods even try to produce re-captioned detailed descriptions by more powerful models rather than existing human-annotated short captions to train their model llavaov ; videollama3 . Therefore, it is urgent to propose a new visual caption benchmark that adapts to modern MLLMs.
@@ 2 Related Work
Visual caption benchmarks. Visual Caption is a fundamental task in computer vision. Early visual caption benchmarks, such as MS-COCO mscoco , NoCaps nocaps , MSR-VTT msrvtt , and VATEX vatex , usually contain a short sentence with limited visual information as the ground truth. They also use metrics like BLEU papinesi2002bleu , CIDER vedantam2015cider , and METEOR banerjee2005meteor to calculate the matching score directly between two sentences, which is easily affected by the sentence style. Recently, from the annotation aspect, DetailCaps detailcaps extracts object-related information from the ground-truth caption, Dream-1K dream1k splits the ground truth and candidates sentences into events. VDC auroracap also extracts the object, background, and camera information from the video captions by question templates. However, they still rely on the ground-truth caption with human-bias, and require existing LLMs to extract and compare multiple times, thus increasing the cumulative error. CompreCap comprecap explores directly annotating the object-related information in image captions, making the benchmarking more interpretable. On the contrary, we are the first time to propose a comprehensive visual caption benchmark covering both image and video data with 6 views and 12 dimensions. For evaluation, most methods only focus on correctness. Dream-1K dream1k and CompreCap comprecap begin to focus on thoroughness and calculate the recall of events or the object coverage in the segmentation map. However, they still remain incomplete as they only evaluate one dimension and limited metrics. We design metrics about both correctness and thoroughness, which may be ignored by previous work. We summarize the comparison with other visual caption benchmarks in Tab. 2, and we are the most holistic on all listed aspects.
@@ 3.1 Multiple Dimension Data Annotation
Dimension design. Motivated by visual generation benchmarks geneval ; vbench ; t2vcompbench , we conclude 6 views and split them into 12 dimensions based on the analysis of caption cases, as shown in Tab. 1. As shown in Fig. 1, we design 9 static dimensions for both video and image, and 4 dynamic dimensions for video, covering most of what makes up a visual caption. We classify dimensions as dynamic or static based on the following principle: descriptions obtainable from a single frame are static, those requiring the entire video are dynamic, which are more related to temporal information. For the object number dimension, the number can be counted statically in an image, and also dynamically in a video, which is more challenging vsibench . We also design the annotation type of each dimension as two types: open-ended, and specific categories. Specifically, we define 9 categories for style, 4 categories for camera angle, and 7 categories for camera movement. The specific categories of each dimension can be found in Fig. 6. See Appendix A.1 for details of dimension design.
@@
Data collection. For convenience and problem simplification, we only collect image data for static dimensions and video data for dynamic dimensions. This is based on the common sense that the video understanding capabilities for MLLMs are usually built upon sufficient image understanding capabilities llavanextvideo ; llavavideo ; videollama3 ; internvl2.5 . Since an image or a video cannot cover all these dimensions of information, we directly collect data for each dimension independently and evaluate each dimension separately. For static dimensions, we mainly collect images from SA-1B sam , COYO-700M coyo700m , Wukong wukong100m , and Wikipaintings wikipaintings , and we also crawl a considerable amount of data from multiple public datasets and websites with CC0 license by ourselves. We also borrow parts of the image data and annotations from CompreCap comprecap for the spatial relation dimension. For dynamic dimensions, we crawl and cut videos for camera movement dimension, borrow videos from Dream-1K dream1k for action and event dimensions, and borrow videos from VSI-Bench vsibench for the dynamic object number dimension. Fig. 5 shows our data sources for each dimension and their proportion.
@@
Pre-Annotation. The annotations may not be unique for different dimensions. For global-related, camera-related, and knowledge-related views, the annotations tend to be unique as an image only belongs to one kind of scene, style, etc. We directly pre-annotate them by extracting the metadata (e.g., style for images in Wikipaintings wikipaintings ), or ask SOTA MLLMs to get a preliminary answer. For object-related, text-related, and temporal-related views, there could be multiple objects, texts, or actions in an image or video. However, it is extremely hard to annotate all objects or actions within an image or a video, as the categories of objects can be divided by almost infinite granularity tang2023visual ; wang2023all ; wang2024all . Therefore, we do not pursue the most comprehensive annotation possible for each single sample, but randomly select only one object from the visual content, and the same for other dimensions, and reflect the accuracy and thoroughness through the evaluation of a large number of samples. We name this strategy as One Represents All. According to the law of large numbers, the distribution of randomly selection can approximate the expectation of covering different granularities of the entire visual content with a large amount of samples, thus ensuring the unbiased nature of the benchmark. Therefore, the key of this annotation strategy is to keep the selection as random as possible. To avoid humans’ bias on selecting, we ask the three SOTA MLLMs, i.e., GPT-4o gpt4o , Gemini-1.5-pro gemini1.5 , and Qwen-VL-Max qwen2vl to list all objects and actions at the granularity they deem appropriate in an image or video, ask PaddleOCR ppocrv3 to list all texts in an image. We finally use Qwen2.5-Max qwen2.5 to merge the results together and randomly select one from the merged list to obtain the pre-annotated results. For further object-related dimensions, e.g., object number, object color, and spatial relation, the object selection follows this strategy, then pre-annotate these attributes by MLLMs.
@@
Data balance. Based on the pre-annotation results for each dimension, we conduct data balance strategy to control the difficulty and diversity. For dimensions with specific categories, we try to make the number of each category similar. For dimensions of open-ended descriptions, we count the frequency of descriptions, suppress the long-tail distribution, keep low-frequency words, ensuring the variety. See Appendix A.3 for examples. The final annotation distribution is shown as Fig. 6.
@@
Data filtering. We finally conduct data filtering to drop harmful visual content and re-balance the data since many of them are modified manually. The final distribution of each dimension is shown in Fig. 6. Some benchmark examples are shown in Appendix A.5.
3.2 Multiple Dimension Evaluation
Caption evaluation. As we collect and annotate the data of each dimension separately, we also evaluate each dimension independently. Different from matching the similarity between the caption and ground-truth sentences, our annotation drops the caption sentence, and we use GPT-4 Turbo gpt4 to take interpretable scores for each dimension. We use a similar prompt template for dimensions with specific categories (i.e., style, camera angle, and camera movement), and use another similar prompt template for other open-set dimensions. See Appendix B.2 for the details of the prompts. Therefore, we can judge the caption into the following three situations: 1) MIS, which means the caption does not mention the corresponding content about the dimension. 2) COR, which means the caption mentions the corresponding content about the dimension, and describes it correctly compared with the annotations. 3) INC, which means the caption mentions the corresponding content about the dimension, but gives a wrong description compared with the annotations. As all data can be judged into these three situations, we can then calculate two metrics: 1) Precision, which represents the accuracy on all samples that the model has mentioned, and thus only considers the correctness. 2) Hit, which represents the accuracy on all annotated samples, no matter whether the dimension is described or missed in the caption, and thus also considers thoroughness. Our Hit is similar to recall, see Appendix A.4 for more discussions. Given the set of all samples as $S(\text{ALL})$, positive samples as $S(\text{COR})$, negative samples as $S(\text{INC})$, missed samples as $S(\text{MIS})$, the metrics can be calculated by:
Precision $=\frac{|S(\text{COR})|}{|S(\text{COR})\cup S(\text{INC})|},$ (1)
Hit $=\frac{|S(\text{COR})|}{|S(\text{ALL})|}.$ (2)
Table 3: The precision and hit of closed-source models and 72B open-source models on all dimensions. The precision represents the accuracy of what the models have described. The hit shows how many visual elements in the image can be described correctly. For video inputs, we send the whole video for Gemini, and uniformly sample 50 frames for GPT due to the API limitation of image number.